What is the probability that a female student will be selected at random from a class
of seven males and three female students, in each of the next two class meetings?
1
Expert's answer
2020-11-27T13:21:18-0500

A probability that a female student will be selected at the meeting is "\frac {3} {10}" .

A probability that a female student will be selected at two meetings is "\frac {3} {10} * \frac {3} {10} = \frac {9} {100}=0.09" or 9%.